[figure: 9.7 cm]

BREAKDOWN DE GENNES SCALING IN Y_{1−x}Dy_{x}Ni_{2}B_{2}C SINGLE CRYSTALS

Z.Drzazga^{a}, A.Winiarska^{a}, D.Kaczorowski^{b} and A.Winiarski^{a} aInstitute of Physics, University of Silesia, ul. Uniwersytecka 4, 40-007 Katowice, Poland

bInstitute of Low Temperature and Structure Resea rch Polish Academy of Sciences P.O. Box 937, Wrocław, Poland

Single crystals of Y_{1−x}Dy_{x}Ni_{2}B_{2}C (1≤ x ≤ 0.7) were grown using arc- melting method. Quality of the crystal was examined by X-ray diffraction and stoichiometry was attested by XPS measurements using electron spectometer PHI 5700/660. Magnetic measurements were carried out with a SQUID magnetometer at the temperature range 1.8 - 300 K in the applied field up 5 T. The superconducting transition temperatures (T_{c}) and Néel temperatures (T_{N}) were determined for investigated samples. Dramatic depressing of superconductivity observed in lightly Y substituted DyNi_{2}B_{2}C leads to a breakdown of the Gennes scaling for T_{c}⟨=T_{N}. Interactions between the rare earth ions and the conduction electrons seem to be very sensitive to the stoichiometry of the constituents.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Z. Drzazga

Address for correspondence : Institute of Physics, University of Silesia ul. Uniwersytecka 4, 40-007 Katowice

Email address : drzazga@us.edu.pl